### v2

# Do Multimodal Large Language Models See Like Humans?

Abstract
11footnotetext: Equal contributions. Project page: https://jiaying.link/HVSBench/
Multimodal Large Language Models (MLLMs) have achieved impressive results on various vision tasks, leveraging recent advancements in large language models. However, a critical question remains unaddressed: do MLLMs perceive visual information similarly to humans? Current benchmarks lack the ability to evaluate MLLMs from this perspective. To address this challenge, we introduce HVSBench, a large-scale benchmark designed to assess the alignment between MLLMs and the human visual system (HVS) on fundamental vision tasks that mirror human vision. HVSBench curated over 85K multimodal samples, spanning 13 categories and 5 fields in HVS, including Prominence, Subitizing, Prioritizing, Free-Viewing, and Searching. Extensive experiments demonstrate the effectiveness of our benchmark in providing a comprehensive evaluation of MLLMs. Specifically, we evaluate 13 MLLMs, revealing that even the best models show significant room for improvement, with most achieving only moderate results. Our experiments reveal that HVSBench presents a new and significant challenge for cutting-edge MLLMs. Diverse human participants attained strong performance, significantly outperforming MLLMs, which further underscores the benchmark’s high quality. We believe that HVSBench will facilitate research on human-aligned and explainable MLLMs, marking a key step in understanding how MLLMs perceive and process visual information.
Figure 1: *
(a) HVSBench
Figure 2: *
(b) Comprehensive comparison.
Figure 3: We are the first to systematically study and assess MLLMs-HVS alignment, introducing diverse HVS-related questions. (a) We propose large-scale and comprehensive HVSBench, with a robust evaluation protocol. (b) We evaluate 13 leading MLLMs with HVSBench across 5 fields, revealing room for improvement and insights for developing HVS-aligned MLLMs. The shown predictions are from GPT-4o. For brevity, only 2 of 10 human scanpaths are shown. Questions and predictions are simplified. Red boxes are NOT in original image.
1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) have shown remarkable progress, achieving impressive performance across diverse vision-language tasks. These include applications such as image captioning [6], visual question answering [19, 36], document understanding [43], mathematic reasoning [35], and more. Such achievements highlight the capabilities of MLLMs in visual perception and vision-language interaction.
Despite the impressive performance of MLLMs on vision tasks, we have limited understanding of why they perform well. It has been demonstrated in previous studies that principles inspired by the HVS and grounded in cognitive science [13], play a vital role in enhancing the performance of backbone models, such as attention-based architectures [51]. However, discrepancies remain in how MLLMs and humans perceive visual information. Human attention is based on innate and learned saliency, while MLLMs often perceive images as arrays of pixel values or feature embeddings. Human visual attention is sequential, adjusting based on context and prior knowledge, while MLLMs process input statically or through fixed-length attention. Human attention can be dynamically influenced and guided by goals, while MLLMs lack the cognition and the ability to “refocus” dynamically, relying purely on trained associations. Given the benefits and importance of HVS-aligned designs, including improving QA and captioning, content generation, downstream tasks, and practical applications, etc., as discussed in Sec. 4.4 , the question of whether and to what extent existing MLLMs align with HVS remains a critical area for research, especially considering the current disparities.
This brings us to a fundamental question: Do MLLMs perceive the world in the same way as humans do? More specifically, do MLLMs fixate on similar regions of interest within an image or follow a similar temporal order as the HVS when perceiving an image? Humans can easily identify objects that capture their attention and perform visual searches based on context. However, in real-world scenarios, MLLMs often perceive visual information differently. For instance, in the first image of Fig. 3, if we ask, “Which object is the most visually prominent?”, most people would choose the center giraffe, while MLLMs identify the right one. Similarly, when searching for a chair in the second image, humans tend to first identify related objects (e.g., a table) and use contextual cues to help locate the chair (e.g., chairs are often near tables), while MLLMs may point to irrelevant areas. However, there has been limited research and challenges on how current MLLMs align with the HVS. On the one hand, existing public vision-language datasets are primarily designed to assess model performance on specific tasks, offering little insight into its alignment with HVS. On the other hand, traditional HVS research has largely focused on low-level, vision-only domains [38], often evaluating models based on masks or heatmaps, making it difficult to assess MLLMs where the primary output is text rather than visual data.
In response to these challenges, we introduce a comprehensive benchmark HVSBench and designed an evaluation protocol suite for MLLMs. Currently, HVSBench contains over 85K questions covering five distinct fields of the HVS, paired with images and answers:
  1. 1.
    Prominence. Test whether the regions MLLMs focus on align with those that are prominent to human perception. Example question in Q1 of Fig. 3.
  2. 2.
    Subitizing. Test whether the number of visually prominent objects for MLLMs matches human perception.
  3. 3.
    Prioritizing. Assess if the order of importance assigned by MLLMs to objects reflects human viewing priorities.
  4. 4.
    Free-viewing. Check if MLLMs can mirror the human attention shift (i.e., sequence of locations that the HSV attends to) in an image during free viewing.
  5. 5.
    Searching. Test if MLLMs can follow a similar sequence of gazes as humans when searching for a specific object in an image. Example in Q2 of Fig. 3.
The questions can be categorized into 13 types based on their phrasing, and the answer types include multiple-choice, counting (i.e., predicting an integer value), sorting, and scanpath prediction. To ensure both quality and variability, we design our benchmark based on a curated collection of large-scale and high-quality datasets [54, 38, 7, 12] focusing on the HVS derived from real-world human studies. For a balanced assessment of each field, we carefully curate and divide the question field so that each field encompasses over 6,707 questions. For the evaluation protocol, while pure exact-match metrics are unreliable due to the limitations of MLLMs in instruction-following and choice labeling, GPT-4-based matching adds bias, high costs, and struggles with complex predictions like scanpaths. Thus, to reduce the matching-caused false-negative and improve evaluation robustness across fields, we design a human-inspired and field-adaptive automatic standardization, taking inspiration from diverse possible predictions and adapting to each field programmatically.
We thoroughly evaluate 13 well-known SOTA MLLMs and human performance on HVSBench, spanning diverse architectures and model scales. This not only provides a direct comparison among these models across multiple aspects of the HVS, but also highlights the significant gap between current MLLMs and humans. Furthermore, our findings reveal critical insights for future improvements: aligning MLLMs with the HVS cannot be achieved merely by incorporating external knowledge of related cues and priors, or by integrating human-generated captions and summaries. Additionally, we stress the value of HVS alignment across domains and their evidences: mimicking human fixation enhances QA/captioning; more intuitive content generation demonstrated by our designed prominence enhancement; improved performance in HVS-specific tasks; and applications in autonomous driving and assistive tools. Our contributions are summarized as follows:
  1. 1.
    We are the first to systematically study the alignment between MLLM and the HVS, pioneering the use of diverse questions across five distinct fields of HVS for model evaluation.
  2. 2.
    We construct HVSBench, a large-scale and comprehensive benchmark with 85,147 multimodal question-answer pairs, which can be used to thoroughly evaluate MLLMs in scenarios that closely mirror the HVS.
  3. 3.
    We propose a robust evaluation protocol with a human-inspired and field-adaptive automatic standardization.
  4. 4.
    We conduct a comprehensive evaluation of 13 popular MLLMs and human using HVSBench. Additionally, we provide new insights and techniques for developing more HVS-aligned and explainable MLLMs, including a prominence enhancement method that shows how HVS-aligned MLLMs improve downstream applications.
2 Related Work
Existing benchmarks for MLLMs focus on assessing their capabilities in understanding and reasoning across modalities. A wide range of benchmarks, such as MMBench [33] and SEEDBench [27], evaluates general multimodal capabilities, including visual perception, reasoning, and comprehension. These benchmarks span diverse tasks, including document understanding [39], visual question answering [37], hallucination detection [31] and mathematical reasoning [35]. While existing benchmarks assess various abilities, they do not fully evaluate the alignment of MLLMs with the HVS. The introduction of HVSBench addresses this gap, focusing on human-centric evaluation to reveal the secret of how MLLMs perceive and process visual information. More related work on different areas (e.g., HVS and MLLM) is in the supplementary material.
Figure 4: Samples of 13 question types in HVSBench with simplified text. Showing GT ranks and scanpath plots for better visuals.
3 HVSBench
In this section, we present our HVSBench in detail. We begin by defining the five distinct fields of HVS, followed by a description of multimodal QA generation, and automatic standardization and evaluation protocol. Fig. 4 shows some examples in our HVSBench. The Sec. 1 and Sec. 4 in supplementary materials provide the rationale behind selecting fields, their corresponding source datasets, additional sample questions and the results for each question type.
3.1 Preliminaries
To evaluate MLLM alignment with the HVS, we construct multimodal QAs based on five key fields, selected to represent diverse aspects of human visual processing and provide a comprehensive assessment. They include:
Prominence. The prominence in HVS enables humans to identify the most visually prominent objects within an image [38], making it a critical application for understanding human visual focus in the HVS.
Subitizing. Subitizing [60] is to quickly and accurately perceive the number of visually prominent objects in a scene. Unlike Prominence, it requires simultaneous attention to multiple items, vital for tasks requiring quick estimation of prominent items’ number, e.g., navigation, crowd analysis.
Prioritizing. Prioritizing in HVS ranks objects within a scene by perceptual saliency [12]. It better captures the dynamic nature of HVS, i.e., the relative visual importance of objects, whereas Prominence and Subitizing focus on static characteristics. This field has broad applications, like autonomous driving, which requires understanding relative saliency for explainable, HVS-driven decisions.
Free-Viewing. Free-viewing is crucial for both behavioral vision in the HVS and computer vision applications. It focuses on the sequence of human gaze behavior in a task-free context [55]. This behavior is bottom-up, driven solely by intrinsic visual processing. As free-viewing closely mirrors human gaze behavior, evaluating a model’s ability to predict both the order and duration of fixations, as well as a full scanpath, reveals its alignment with the HVS.
Searching. Searching [54] focuses on human gaze behavior in task-driven contexts, such as object search, where attention is top-down and influenced by contextual information, like object context and semantic relationships. Unlike free-viewing, searching enhances human efficiency and flexibility [54]. Therefore, aligning MLLMs with the search domain may lead to similar improvements.
The selection of the five key fields are grounded in the dual-process theory introduced in [3] and grounded by cognitive science in the introduction section of [10] of HVS, which distinguishes between bottom-up (stimulus-driven) and top-down (goal-directed) processes. Bottom-up processes mainly include prominence [20], free-viewing [21] and subitizing [50]. Top-down processes mainly include prioritizing [17] and searching [15]. These fields were chosen because they span the majority of human visual behavior, as shown in foundational reviews on visual scene processing [18], which emphasizes saliency (Prominence), attention dynamics (Free-Viewing), and search (Searching) as critical components of HVS. While others may exist, these five fields are the most widely studied and theoretically validated in HVS to the best of our knowledge.
3.2 Benchmark Curation
Based on the above definitions, we design an automatic multimodal QA generation paradigm to convert source annotations into different forms of QAs for MLLM evaluation.
Multimodal QA Generation. Since the annotations in the selected datasets are not in the multiple-choice QA format, we automatically transform the ground-truth annotations into this format automatically with human verification. Specifically, we manually created 13 types of question templates. Following previous practices [33], most of our evaluations are conducted using multiple-choice QAs rather than open-ended ones, except for fields with well-defined outputs such as reporting a positive integer (e.g., subitizing) or a list of coordinates (e.g., scanpath prediction). These formats align directly with their problem metrics, enabling precise and unbiased evaluation. Open-ended answers, by contrast, often require scoring by LLMs or user studies, which can introduce evaluation bias or require manual intervention based on previous practices [33]. Then, we create the corresponding answer options as follows.
Variations in perception.We leverages GT human data in peer-reviewed cognitive science datasets [7, 12, 38, 54] rigorously designed for diversity and bias control. For example, Fig. 3’s source [12] employs 8 annotators (a common sample size [30] in saliency research) in diverse culture (EU/CN), race (Black/White/Asian), and gender (4M/4F), collected by strict protocols (e.g., eye-tracking monitor) to minimize bias (Sec. 3 of [12]). This ensures our benchmark captures shared human attention tendencies, aligning with common practices in HVS research.
(1) Template-Based Construction: For most questions, we generate answer options directly from GT human annotations. For example, in fields like Prominence, the options are derived from bounding boxes or other directly annotated features. To ensure diversity in the questions, we design multiple templates for each question type, following the methodology outlined in [32]. These templates provide variation while maintaining the focus of the field. For instance, in Prominence, we use templates such as:
  • Which object is the most salient in the image?
  • Which object is the most visually prominent in the image?
  • Which object attracts the highest visual attention?
All templates are manually created by humans to ensure accuracy and relevance.
(2) LLM-Based Refinement: Note that some MLLMs may not be able to utilize the coordinate data well, but they are more suitable for interpreting in natural language. We design a bunch of questions that do not explicitly involve numerical data. However, such data cannot be easily obtained due to the limited original label types of data sources. For scenarios where bounding box coordinates are not explicitly required and the field involves natural language references (e.g., Between these two objects — [A] $<$obj1$>$ [B] $<$obj2$>$ — please select the option representing the more salient object.), we use a large language model (LLM) [1] to generate description in natural language for the options ([A] the person on the left side; [B] the bicycle on the right side). Specifically, we use GPT-4 [1] to describe objects based on their bounding boxes and to create plausible but false options from non-salient objects. This method ensures that the QA matches real-world applications where natural language complements visual data.
(3) Human verficication. To address potential ambiguities introduced by LLMs or the complexity of image content in (2), human verification is employed to ensure that the referred objects in natural language descriptions correspond accurately to the intended targets. This additional step improves the accuracy and reliability of generated QAs.
Answer Option Processing. For multiple-choice questions, we randomly sample answer options from the available candidates and shuffle their order. This approach enhances robustness by minimizing biases arising from fixed option orders or repetitive patterns in the answer choices. For scanpath-related questions (i.e., free-viewing and searching), following [54], we randomly sample human scanpaths from different images. This ensures that the alternative answer options are also derived from human data, making them more natural. By avoiding reliance on computer-generated scanpaths, we ensure that the evaluation remains unbiased and fair for the evaluated LLMs.
To this end, we produced 85K multimodal QAs based on annotations from 71K images, covering 13 question types across 5 key fields in HVS, including prominence, subitizing, prioritizing, free-viewing and searching. Table 1 and Table 2 show statistics of our HVSBench, which is large and diverse, covering various answer types and tasks.
  Prominence Q1 Subitizing Q2 Prioritizing Q3-Q7 Free-Viewing Q8-Q10 Searching Q11-Q13
Total QAs 85,147 8,389 (11%) 18,105 (24%) 26,309 (34%) 17,090 (22%) 6,707 (8%)
Question Categories Average # of Obj. Max./Avg. Q Length Max./Avg. A Length Max./Avg. # of Choice
13 3.2 104/45.8 5/1.4 17/3.9
Table 1: Key statistics of HVSBench. Fixation prediction statistics are excluded due to lengthy answers.
  Benchmarks # Total QA Answer Type Evaluation Purpose
       General Propose Benchmarks
MME [16] 1457 Y/N General
MMBench [33] 3,217 A/B/C/D General
SEED-Bench [27] 19,242 A/B/C/D General
MVBench [29] 4,000 A/B/C/D General
       Special Propose Benchmarks
POPE [31] 3,000 Y/N Object Hallucination
MathVista [35] 6,141 Choice Free-form Math Reasoning
GlitchBench [42] 593 Free-form Detect Game Glitch
  HVSBench (Ours) 85,147 Choice Integer Scanpath Alignment with HVS
Table 2: Comparison to other benchmarks.
3.3 Automatic Standardization and Evaluation
Automatic Standardization. In evaluation protocols, pure exact-match metrics prove unreliable due to the limitations of MLLMs in instruction-following. For example, in Fig. 5, even when we explicitly specify the output format and provide example outputs for reference, MLLMs still produce predictions with inconsistent and somewhat random formatting. Meanwhile, the widely adopted LLM-based matching approaches [14] introduce biases, incur high costs due to multiple evaluation passes, and struggle with complex predictions such as scanpaths. To address these challenges, we propose a human-inspired, field-adaptive automatic standardization method. It minimizes false negatives caused by matching errors and enhances evaluation robustness by programmatically adapting to diverse prediction formats across the five fields in our benchmark.
To achieve automatic standardization programmatically, we adopted a human-inspired, field-adaptive approach. Specifically, we observed that the responses to each question type are not open-ended but instead follow discernible patterns. Building on this insight, for each field, we randomly sampled $10$ example questions for each question type. On these example questions, we then collected multiple rounds of random predictions from all models evaluated in our experiments, as well as predictions from over $20$ different human participants and their human-annotated standardizations. With these diverse and complex predictions across fields, we programmatically defined an automated standardization process. This process ensures robust performance without introducing errors, even when applied to predictions outside the seen predictions.
For the Prominence field, the automatic standardization cleans and normalizes predictions into consistent main choice labels (e.g., ‘A’, ‘C’, etc.) from various input formats, using regex to detect explicit or implied labels in diverse phrasing. For Subitizing, the automatic standardization processes predictions to extract integer values, including direct numbers, text-based numbers (e.g., “three”), or implied quantities. When no valid number is found, it defaults to the average ground truth value in 3.2116. For Prioritizing, except for choice answers that are processed the same as in the Prominence field, our automatic standardization detects and normalizes answers of sequence (sorting) via order-related phrasing through regex patterns. In Free-viewing and Searching fields, The automatic standardization involves extracting and normalizing scanpath data (X, Y for coordinates and T for durations) from diverse input formats, including text, JSON, and various phrasing styles. It also addresses irregularities like mismatched list lengths and incomplete scanpaths. Refer to Sec. 2 in supplementary for predictions and their corresponding standardized outputs in each field. We illustrate the superiority of our automatic standardization over LLM-based matching in Fig. 5. In our experiments, GPT-4 was prompted with the full question, the full prediction, and the text prompt: “Please extract the prediction with the correct format.” Refer to Sec. 3 in supplementary for detailed inputs, experimental settings, and pseudo-code for our automatic standardization.
Figure 5: Illustration of our automatic standardization, which robustly formats predictions without introducing errors. In contrast, LLM-based matching (e.g., GPT-4) is both costly and prone to errors, such as failing to extract the correct format or introducing biases by predicting unrelated outputs.
Evaluation Metrics. In the Prominence field in which are choice questions, the accuracy can be calculated in $\text{Acc}=\frac{\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i})}{N}$, where $y_{i}$ is the correct answer for the $i$-th question, $\hat{y}_{i}$ is the prediction for the $i$-th question., $\mathbb{I}(\hat{y}_{i}=y_{i})$ is the indicator function that equals 1 if $\hat{y}_{i}=y_{i}$ and 0 otherwise, and $N$ is total number of predictions. For the Subitizing field, we adopt widely used metrics, including Mean Absolute Error (MAE), Root Mean Square Error (RMSE), as well as Exact Match Accuracy (Acc). For Prioritizing field, the answer types include single-choice and sequence (sorting). The single choices are evaluated via accuracy. To evaluate whether the predicted order of prioritized items matches the ground truth exactly, Exact Match Accuracy (Acc) is also used. In Free-viewing and Searching fields, we employ the widely adaptive MultiMatch [21] similarity to evaluate generated scanpaths. It measures scanpath similarity across multiple dimensions: shape, direction, length, and position. In our main results, we focus on the direction (”M-Dir”) and position (”M-Pos”) dimensions for simplicity. Following standard procedures [54], predictions are cropped to a maximum length of 6, and the final scores with each prediction being the mean of the MultiMatch calculated across all 10 GT scanpaths in the image.
[…]
Q1 Q2 Q2 Q2 Q3-Q7 Q8,Q10 Q9 Q9 Q11,Q13 Q12 Q12
[…]
       Human Performance
Human 0.7438 0.7683 0.7143 0.6349 1.3333 0.7025 0.7560 0.6963 0.8863 0.9000 0.7617 0.9181
Baselines
Random guess 0.2806 0.1994 0.1909 1.9996 2.4336 0.2949 0.3750 0.6079 0.8304 0.3752 0.5425 0.8160
Proprietary MLLMs
GPT4-o [1] 0.3946 0.3139 0.4512 1.3445 3.2614 0.3621 0.3737 0.5917 0.8042 0.4015 0.5106 0.7834
Gemini-1.5-Flash [45] 0.3886 0.3323 0.5106 1.3070 3.0978 0.3283 0.3804 0.6128 0.8392 0.4040 0.5199 0.8206
OpenSource MLLMs
MiniGPT4-v2 [62] 0.1545 0.0879 0.1029 2.4939 3.0657 0.1810 0.1830 0.5017 0.4990 0.2294 0.5364 0.5051
DeepSeek-VL [34] 0.3655 0.3223 0.4544 1.2471 2.2514 0.3327 0.3445 0.5118 0.5450 0.3516 0.4790 0.7130
Idefics [26] 0.2488 0.1905 0.2386 2.1971 3.1163 0.2280 0.3230 0.5422 0.6236 0.2494 0.5293 0.4895
Idefics2 [26] 0.3067 0.2015 0.2990 2.2078 3.8573 0.2971 0.3637 0.5151 0.5525 0.3865 0.5322 0.7619
Idefics3 [25] 0.3552 0.2149 0.4852 1.1055 1.9567 0.3272 0.3411 0.5329 0.6310 0.2843 0.5407 0.6856
LLaVA-Next [32] 0.3460 0.3223 0.3961 1.3727 2.2675 0.3341 0.3274 0.5249 0.5998 0.3192 0.5083 0.7883
LLaVA-OneVision [28] 0.4035 0.4640 0.4517 1.1206 1.9671 0.3495 0.3888 0.5490 0.8177 0.4913 0.5047 0.7295
mPLUG-Owl3 [58] 0.3076 0.3309 0.2688 2.8941 4.6162 0.3305 0.3117 0.5028 0.4941 0.2668 0.5368 0.5592
Qwen2-VL [52] 0.4076 0.4103 0.5090 1.4698 2.6747 0.3901 0.3182 0.5199 0.7178 0.4214 0.4610 0.8200
InternVL2.0 [8] 0.3082 0.3101 0.3405 2.3458 3.6068 0.3405 0.3409 0.5028 0.7090 0.2968 0.5016 0.7282
MiniCPM-V 2.6 [57] 0.3476 0.3748 0.4491 1.2471 2.1137 0.2989 0.3023 0.5132 0.5589 0.3416 0.4953 0.7107
Table 3: HVSBench Leaderboard. The results of 13 leading MLLMs reveal significant room for improvement.
4 Experiments
4.1 Evaluation Details
We use different metrics for extensive evaluations. In the Subitizing, the accuracy is computed using random sampling based on the frequency distribution of subitized object counts. For MAE and RMSE, the average number of salient objects across all samples is used as the predicted value. For MultiMatch metrics (i.e., M-Dir and M-Pos), following [54], a random guess sample is a randomly selected human scanpath of a different image. Our human evaluation involves 10 participants from diverse backgrounds considering individual differences [23]. For MultiMatch, we use a commonly used oracle method [54, 41], which compares each subject’s scanpath as a prediction to others’ as ground truth and averages the results.
We select recent open-source MLLMs for validation, including MiniGPT4-v2 [62], DeepSeek-VL [34], Idefics-series [26, 25] LLaVA-Next [32], LLaVA-OneVision (LLaVA-OV) [28], mPLUG-Owl3 [58], Qwen2-VL [52], InternVL2.0 [8] and MiniCPM-V 2.6 [57]. We utilize the $\sim$7B-scale versions of these models, a widely adopted choice for balancing efficiency and effectiveness. We also include GPT4-o [1] and Gemini-1.5 Flash [45], two representative proprietary MLLMs, as key benchmarks in our evaluation. All experiments are conducted using VLMEvalKit [14] on the same platform for consistency and fairness. We allocate 10% of HVSBench for evaluation and reserve the remaining 90% for further explorations, like instruction-tuning MLLMs.
Figure 6: Qualitative results. Note that the bounding boxes, the scanpaths and the GT labels from the source datasets (e.g., GT saliency rank instances) are for visual clarity and are not used in the input images for evaluation. Text is partially omitted due to space limitations.
4.2 Main Results on HVSBench
Quantitative Evaluation. Table 3 presents the evaluation results on HVSBench, highlighting that current open-source and proprietary MLLMs still underperform in aligning with HVS. Diverse human participants achieved strong performance (OA: 0.74). Given individual differences in saliency and cognitive science [23], this performance is both high and reasonable, which further validates the quality of our benchmark. Humans clearly outperform both random guessing (OA: 0.28) and MLLMs (best: 0.41) in all metrics. It is worth noting that MiniGPT4-v2, an early MLLM, has a bad performance, even worse than a random guess. This is because it often generates invalid outputs (refusing to answer or irrelevant responses) instead of outputting an option. Taking the outperforming MLLMs as a different kind of examples, GPT4-o falls short in several metrics, such as Prominence, Prioritizing, and Searching accuracy, where it performs worse than several open-source models such as Qwen2 [52]. Qwen2 [52] achieves the best overall performance, with LLaVA-OV [28] following closely behind. Both models even surpass GPT4-o overall, demonstrating the potential of recent open-source MLLMs.
In Free-viewing, only LLaVA-OV [28] and Gemini-1.5-Flash [45] outperform random guessing in terms of accuracy. The situation improves slightly in the field of Searching. Specifically, both LLaVA-OV [28] and Qwen2 [52], and two proprietary MLLMs, greatly outperform random guessing in accuracy. A possible explanation is that in Free-viewing, which is a pure vision field that reflects HVS behavior without conditioning, these models exhibit behaviors that differ from humans. In contrast, Searching involves more defined objectives and patterns that are uniform and predictable, making them easier for MLLMs to align with.
However, almost all methods perform poorly on scanpath similarity metrics (“M-Dir” and “M-Pos” in Free-viewing (Q9) and Searching (Q12)), underperforming random guessing, except for Qwen2 [52] in searching. These results suggest significant room for scanpath prediction for current MLLMs. The Sec. 6 in supplementary materials shows the detailed benchmark on all 13 question types.
Benchmark Examples and Predictions. Fig. 6 shows benchmark examples and qualitative results of the three most representative models that performed best on our benchmark. We include the ground truth saliency information based on natural viewing by human observers for reference. The top-left illustrates a sample from the Prominence field. A human can easily identify the person closest to the center as the most salient. However, both GPT-4o and Qwen2 [52] fail in this case. The top-right corner shows a sample from the Subitizing field. Here, while GPT-4o and LLaVA-OV [28] fail to predict the correct quantity, Qwen2 [52]’s prediction can be automatically standardized to match the correct result. The bottom-left corner features a sample from the Prioritizing field, where all three models fail. The bottom-right corner displays a sample from the Free-Viewing field. Option A reflects a clear pattern of human attention, while the human scanpath shown in option B is unrelated to the image. Despite this, all models incorrectly selected B. The bottom-right shows a hard case in our HVSBench where MLLMs struggle to choose the human free-viewing scanpath, highlighting that their behavior in purely vision tasks differs from HVS. This may be because they are optimized for object recognition, not for simulating human-like viewing behaviors. Refer to Sec. D in the supplementary for more qualitative results.
MLLMs’ Explanation on its Choice. To understand why GPT-4o errs on the prioritization task (bottom left) and to identify its focus of attention, we asked GPT-4o to explain its reasoning: The van is large and occupies a significant portion of the image. Its size and central positioning make it visually dominant. The person near the van is secondary in prominence, as they are close to the van but are relatively smaller in size. The person standing on the right is smaller and less centrally located, reducing its visual dominance. The person on the far right is further isolated, diminishing their salience. Based on size, position, and visual weight, the van appears to be the most salient object in this image. It highlights an over-reliance on physical attributes like size and position, overly associating these with salience. It also reveals a limitation in GPT-4o’s attention mechanism: neglecting social or human-centric cues, which might prioritize smaller yet contextually significant objects or people.
4.3 Ablation Study
Model Size. We evaluate the impact of model size on the performance in HVSBench by testing models with different parameter counts. For this study, we select two representative methods: DeepSeek-VL [34] and GPT-4o [1], from both open-source and proprietary MLLMs, to provide a comprehensive analysis. As shown in Table 4, larger MLLMs generally outperform smaller ones across all metrics. It suggests that increasing model size leads to better alignment with HVS for MLLMs. Refer to the supplementary for more experiments on this aspect.
[…]
  GPT4-o mini [1] N/A 0.3126 0.4480 0.3312 0.3560 0.3766
GPT4-o [1] N/A 0.3139 0.4512 0.3621 0.3737 0.4015
DeepSeek-VL [34] 1.3B 0.1758 0.2513 0.2950 0.3188 0.2843
7B 0.3223 0.4544 0.3327 0.3445 0.3516
Table 4: Ablation study of The number of params. PO, SU, PI, FV and SE means “Prominence”, “Subitizing”, “Prioritizing”, “Free-viewing”, “Searching”, respectively.
Human Captions or Descriptions Improve Alignment? It is possible that certain annotations, like human captions from COCO caption and detailed descriptions from LLaVA-Instruct-150K, implicitly reflect the HVS. Therefore, we evaluated whether adding detailed descriptions (“Detail.”) or short captions (“Cap.”) could improve the performance of the base GPT-4o model.
As shown in Table 5, in the field of Prominence and Subitizing, adding detailed descriptions and captions improved performance, suggesting that additional information helps the model better identify and subitize visually dominant elements. In Prioritizing, while detailed descriptions provided some improvements, captioned input worsened the score, suggesting that brief caption information may not be sufficient for understanding object importance in a human-aligned way. In Free-viewing, both detailed descriptions and captions actually lowered performance. These results suggest that neither detailed descriptions nor captions can effectively capture the natural attention shifts humans exhibit in free-viewing scenarios. In Searching, captions slightly improve accuracy, while detailed descriptions lead to a decrease. One possible reason is that humans typically shift their attention rapidly between the most salient objects during searching. Captions, by prioritizing only the most salient features, might better mirror this fast-paced, targeted attention. Detailed descriptions, however, encourage the model to consider a broader range of objects, leading to attention that doesn’t align with humans during searching.
These results suggest that providing more context through human captions or detailed descriptions can lead to better performance in some evaluation criteria, aligning MLLMs with the HVS cannot be achieved merely by integrating human-generated captions and summaries, especially in Prioritizing, Free-viewing and Searching fields.
[…]
  GPT-4o [1] 0.3139 0.4512 0.3621 0.3737 0.4015
Detail. 0.4274 0.4809 0.3799 0.3465 0.3791
Cap. 0.4188 0.4852 0.3464 0.3605 0.4190
Table 5: Ablation study of question prompt.
Field-Specific Hints Improve Alignment? Since definitions and prior knowledge of the fields also provide cues for the HVS, we examine whether adding Field-Specific Hints can enhance model performance. For example, the simplified hint for the Prominence field is: The detection of salient objects aims to simulate the human visual perception system by identifying and localizing the most visually striking object(s) in a scene. … (omitted), cues such as color contrast, spatial bias, and depth contrast also influence saliency. Refer to Sec. E in supplementary for full hints in each field.
Table 6 summarizes the results on GPT-4o. Our results show an improvement in the Prominence, Subitizing, and Prioritizing fields but a noticeable decrease in performance for Free-viewing and Searching. This decline suggests that too much contextual information may hinder the model’s focus on the raw visual features necessary for free-viewing tasks. A possible explanation is that prior knowledge might cause the model to focus immediately on specific parts of the image, disrupting the natural temporal sequence of the human gaze. This could lead to fixation sequences that are less representative of human free-viewing or searching.
[…]
  GPT-4o [1] 0.3139 0.4512 0.3621 0.3737 0.4015
0.4212 0.5005 0.3803 0.3476 0.3342
Table 6: Ablation study of system prompt.
4.4 Discussion: Benefits and Applications
While human perception mimicry may not be universally beneficial, but we stress its value in key tasks. (1) Mimicking human perception can significantly improve tasks requiring precise visual grounding, such as QA and captioning. In table 4 of [53], models mimicking human fixation mechanisms achieve shorter search lengths in QA tasks, improving efficiency and accuracy. (2) Content generation models better aligned with the HVS can produce more reasonable outputs. Take the prominent field for example, we design a Cropping-Based Prominence Enhancement to illustrate. Specifically, we examine how GPT-4o crops the image to enhance the prominence of one object: a photo. GPT-4o with a task-specific hint generates a reasonable analysis and successfully crops the image to highlight the photo, compared to the result without hint, demonstrating better alignment with HVS. This can be directly applied to automated design, context-aware content generation, and visual storytelling.
Figure 7: One generation application: Prominence Enhancement.
(3) In HVS-related downstream tasks, models better aligned with HVS perform better. For example, Qwen2-VL outperforms MiniCPM-V in our HVSBench and also in general tasks such as human perception, visual illusion (MMMU, HallusionBench) , but is comparable or worse in tasks unrelated to HVS like math reasoning (MathVision). (4) Practical applications like autonomous driving, perception in robots, assistive tools for the visually impaired, and tasks directly related to the HVS (e.g., saliency ranking), further demonstrate the benefits where HVS alignment ensures intuitive, user-centered outputs. Based on these four aspects, we believe mimicking human perception will benefit machine vision.
5 Conclusion
In this paper, we explore the alignment between MLLMs and the human visual system. We introduce HVSBench, a novel large-scale benchmark designed to evaluate MLLMs on vision tasks that closely mirror human perception. It consists of 85K multimodal QAs across 13 categories and 5 fields, accompanied by a robust evaluation protocol. Our experiments demonstrate that HVSBench poses a new, significant challenge for state-of-the-art MLLMs, highlighting considerable room for improvement. We believe HVSBench will drive the development of more human-aligned and explainable MLLMs, offering critical insights into how these models perceive and process visual information.
[…]
Supplementary Material
The supplementary materials include:
  1. 1.
    Section A: Preliminary Settings for HVBench
    This section explains the rationale behind the selection of fields for HVBench. It also provides details about the corresponding source datasets used for each selected field.
  2. 2.
    Section B: Sample Prediction and Standardization
    We present the predictions generated by the models and the corresponding outputs after applying our standardization process for each field. This ensures consistency and comparability across different fields.
  3. 3.
    Section C: Detailed Settings
    This section provides detailed settings in our experiments and standardization pipeline, including the inputs used in our experiments, the detailed experimental settings, and the pseudo-code for our automated standardization pipeline.
  4. 4.
    Section D: Additional Qualitative Examples and Results
    We include extra qualitative examples showcasing sample questions and the predictions made by MLLMs under each field.
  5. 5.
    Section E: Field-Specific Hints in Each Field
    For each field, we provide specific hints or context that help understand tasks.
  6. 6.
    Section F: Detailed Benchmark Results and Disucussions
    This section contains a detailed report of benchmark results for all 13 question types, offering a comprehensive view of the performance across various fields and question types. We also provide additional ablation studies and related discussions for them.
  7. 7.
    Section G: More Related Work
    We include more related work for different research areas to showcase the necessity of our HVSBench.
Appendix A Preliminary setting for HVBench
To evaluate the extent to which multimodal large language models (MLLMs) align with the HVS, we construct multimodal queries based on five fields designed for different datasets. These fields are critical for assessing whether MLLMs perceive and interpret visual information in a manner akin to humans. For a fair and comprehensive evaluation, the five fields are carefully curated to capture different aspects of the HVS. These fields include:
Prominence. The prominence in HVS enables humans to identify the most visually prominent objects within an image [38], making it a critical application for understanding human visual focus in the HVS. In this field, we choose the SIFR dataset [12] to construct our benchmark data. SIFR is a dataset for relative saliency ranking consisting of 8389 images with 52,173 annotated instances.
Subitizing. Subitizing [60] is to quickly and accurately perceive the number of visually prominent objects in a scene. Compared to Prominence, it requires simultaneous attention to multiple elements. It is crucial in real-world scenarios where humans need to quickly estimate the number of prominent items, such as in navigation or crowd analysis, facilitating fast decision-making in tasks like navigation, searching, and choice-making in the HVS. We choose SIFR dataset [12] and SIS10K [38] for this field since the original subitizing dataset [60] is no longer available. SIS10K is a large-scale salient instance segmentation dataset. SIS10K comprises 10,300 images with meticulously annotated instance-level bounding boxes and masks, surpassing the earlier binary-masked datasets. Unlike traditional datasets that often fail to provide instance-level annotations, SIS10K enables the development of instruction-based data for multimodal QA systems. As suggested in the relevant work, datasets with instance-level salient object annotations are ideal for this field. However, binary salient object detection datasets do not provide instance-level labels, which are critical for accurately quantifying the number of salient objects present in the input data. This limitation highlights the importance of using datasets that explicitly support instance-level annotations to ensure reliable performance in subitizing. Without such data, models may struggle to distinguish between individual salient objects, particularly in complex scenes with multiple or overlapping objects.
Prioritizing. Prioritizing in HVS enables humans to rank objects within a scene based on their perceptual saliency [12]. It better captures the dynamic nature of HVS, i.e., the relative visual importance of objects, whereas Prominence and Subitizing focus on static characteristics. This field has broad applications, like autonomous driving, where understanding relative saliency is essential for explainable, HVS-driven decision-making. In this field, we choose the SIFR dataset [12] to construct our benchmark data. Unlike other ranking datasets, the salient instances in SIFR were determined based on clustering and thresholding on real-world human fixation, ensuring a better alignment with the saliency rank and the real attention model in the HVS.
Free-Viewing. Free-viewing is an important behavior of HVS. Free-viewing (bottom-up) gaze path prediction [55] focuses on modeling and forecasting human gaze behavior in a task-free context, driven solely by the intrinsic saliency of visual stimuli. This involves predicting where humans are likely to fixate on an image based on visual properties such as color, contrast, and texture, rather than external goals or instructions. In our HVSBench, we utilize the COCO-FreeView dataset[7] to construct the assessment data for this field. COCO-FreeView [7] is a dataset containing 6202 images with about 300,000 fixations viewed by human subjects under a free-viewing condition without specific search goals. Each image is annotated with fixation points represented by their coordinates ($x,y$) and the duration of gaze (time $t$) at each fixation. This dataset is particularly valuable for understanding the dynamics of bottom-up attention mechanisms as it reflects human visual exploration in a naturalistic and unbiased setting. By incorporating such data into our benchmark, we aim to rigorously evaluate the accuracy and interpretability of attention models in replicating human-like scanpaths and understanding the intrinsic properties that guide gaze allocation in free-viewing scenarios.
Searching. Searching [54] focuses on human gaze behavior in task-driven contexts, such as object search, where attention is top-down and influenced by contextual information, like object context and semantic relationships. Unlike free-viewing, searching enhances human efficiency and flexibility [54]. Therefore, aligning MLLMs with the search domain may lead to similar improvements. In this field, we employ COCO-Search18 dataset [54] as our primary dataset. COCO-Search18 is the largest high-quality dataset for goal-directed attention, specifically designed to capture human fixation behaviors during visual search tasks. It includes 6202 images annotated with nearly 300,000 goal-directed fixations from 10 participants, each searching for one of 18 target-object categories. We use the standard target-present split. By leveraging COCO-Search18, our framework can rigorously assess how well models replicate human scanpaths and predict task-driven attention allocation. This dataset is crucial for advancing computational models of goal-directed attention, bridging gaps between human and machine visual systems, and enabling practical applications such as robotic vision and human-computer interaction.
Appendix B Sample Prediction and Standardization
Sample Prediction Standardization
“X = [0.49, 0.57, 0.56, …] Y = [0.53, 0.53, 0.51, …] T = [316, 148, 123, …]” X=[0.49, 0.57, 0.56], Y=[0.53, 0.53, 0.51], T=[316, 148, 123]
“X = [0.49, 0.57, 0.56, 0.75, 0.85, 0.95]” X=[0.5], Y=[0.5], T=[0]
“the result scanpath is X = [0.49, 0.57, 0.56], Y = [0.53, 0.53, 0.51], and T = [316, 148, 123]” X=[0.49, 0.57, 0.56], Y=[0.53, 0.53, 0.51], T=[316, 148, 123]
[…]
Table 7: Standardized predictions for Q9, Q12 in Free-viewing and Searching fields.
Table 7 shows the standardized predictions for Q9, Q12 in the Free-viewing and Searching fields.
Sample Prediction Standardization
10 10
“22” 22
“there is 1 sample” 1
“three” 3
“Five” 5
“8.” 8
“num of people in this image is 5” 5
“B” GT_avg
“There are five salient objects in the image: four boats and one flower.” 5
Table 8: Standardized predictions for Q2 in the Subitizing field.
Table 8 shows the standardized predictions for Q2 in the Subitizing field.
Sample Prediction Standardization
“[A]” A
“A.” A
“So the result is B” B
“A clock is a clock. So the answer is B” B
Table 9: Standardized predictions for Q1, Q4, Q5, Q6, Q7, Q8, Q10, Q11, Q13 in the Prominence, Prioritizing, Free-viewing, and Searching fields.
Table 9 shows the standardized predictions for Q1, Q4, Q5, Q6, Q7, Q8, Q10, Q11, Q13 in the Prominence, Prioritizing, Free-viewing, and Searching fields.
Sample Prediction Standardization
“[A]” A
“A.” A
“D B C A” DBCA
“the answer is only B” B
“its A and B” AB
“[A] and [C]” AC
“The order between the A,B,C should be B$>$C$>$A” BCA
“Between A,B,C, so the answer is BCA” BCA
Table 10: Standardized predictions for Q3 in the Prioritizing field.
Table 10 shows the standardized predictions for Q3 in the Prioritizing field.
Appendix C Detailed Settings
Detailed inputs and experimental settings. To clearly illustrate the evaluation settings used in HVSBench and described in our main paper, we provide the full question prompt below.
For the top sample in Fig. 3 in the main paper, the full question is:
“[Requirement] Predict eye fixation points for searching for a sink in the provided image. Output the fixation points as three separate lists, each containing 6 values. The lists should represent – X = The X-coordinates of the fixation points, length 6, normalized relative to the image width, formatted to two decimal places (range: 0.00 to 1.00). Y = The Y-coordinates of the fixation points, length 6, normalized relative to the image height, formatted to two decimal places (range: 0.00 to 1.00). T = The fixation duration at each point in milliseconds, length 6, sum not exceeding 5 seconds, formatted as integers. [Example Output for Another Image] X = [0.50, 0.48, 0.44, 0.66, 0.67, 0.50] Y = [0.49, 0.45, 0.37, 0.26, 0.20, 0.50] T = [73, 193, 95, 635, 592, 312]”
For the middle sample in Fig. 3 in the main paper, the full question is:
“[Requirement] Predict eye fixation points for searching for a tv in the provided image. Output the fixation points as three separate lists, each containing 6 values. The lists should represent – X = The X-coordinates of the fixation points, length 6, normalized relative to the image width, formatted to two decimal places (range: 0.00 to 1.00). Y = The Y-coordinates of the fixation points, length 6, normalized relative to the image height, formatted to two decimal places (range: 0.00 to 1.00). T = The fixation duration at each point in milliseconds, length 6, sum not exceeding 5 seconds, formatted as integers. [Example Output for Another Image] X = [0.50, 0.48, 0.44, 0.66, 0.67, 0.50] Y = [0.49, 0.45, 0.37, 0.26, 0.20, 0.50] T = [73, 193, 95, 635, 592, 312]”.
The full prediction is:
“To extract the fixation points from the provided image, I will use the following steps: 1. Load the image and convert it to grayscale. 2. Apply a Gaussian blur to the image to reduce noise and smooth out the features. 3. Use the Eye Fixation Algorithm to detect the fixation points in the image. 4. Extract the fixation points’ coordinates (X and Y) and durations (T) from the output of the algorithm. Here’s how I implemented the algorithm in Python using OpenCV and scikit-image libraries: python”
import cv2
from skimage import io, filters
# Load the image
img = io.imread(’image.jpg’, as_gray=True)
# Apply Gaussian blur to reduce noise
blurred = filters.gaussian_filter(img, sigma=5)
# Detect fixation points using the Eye Fixation Algorithm
fixations = filters.eye_fixation(blurred, threshold=0.5, size=(10, 10))
# Extract fixation points’ coordinates and durations
X = []
Y = []
T = []
Our input for LLM extraction is:
f“Question: “{question}” Prediction: “{pred}” Please extract the prediction with the correct format.”.
And the full LLM Extraction is:
“The provided response doesn’t complete the prediction in the required format. I will generate the prediction based on the specified format. Prediction Output: X: [0.52, 0.46, 0.39, 0.61, 0.68, 0.50] Y: [0.47, 0.42, 0.34, 0.28, 0.22, 0.51] T: [84, 176, 112, 824, 604, 300] This output meets the following requirements: X and Y are normalized to two decimal places (range 0.00 to 1.00). T represents fixation durations in milliseconds, with the sum not exceeding 5000 ms (5 seconds).”
Pseudocode for our standardization.
Algorithm standardization_Q9Q12:
Input: list_input_scanpaths (list of scanpath strings)
Output: cleaned_scanpaths (list of parsed scanpaths as dictionaries)
1. Initialize an empty list cleaned_scanpaths.
2. For each scanpath in list_input_scanpaths:
2.1 Initialize scanpath_dict with default values:
scanpath_dict = {“X: [0.5], Y: [0.5], T: [0]}
2.2 If scanpath contains ”“‘json”:
a. Extract JSON content using string operations.
b. Parse the JSON content to extract X, Y, and T values.
c. Update scanpath_dict with parsed values, if valid.
2.3 Otherwise, use regular expressions to find matches for:
a. X-coordinates
b. Y-coordinates
c. T (time or duration)
Use helper function parse_values to clean and convert matched strings into lists.
2.4 Update scanpath_dict with parsed X, Y, and T values.
2.5 If X, Y, and T lists are not of equal length:
Reset scanpath_dict to default values.
2.6 Append scanpath_dict to cleaned_scanpaths.
3. Return cleaned_scanpaths.
Helper function parse_values(match):
Input: match (regular expression match object)
Output: list of numeric values
1. If match exists:
a. Remove invalid characters (e.g., “…) and split the string by commas.
b. Convert valid strings into float or int values.
c. Return the cleaned list of values.
2. Otherwise, return an empty list.
Figure 8: Pseudo Code
Pseudo-code 8 shows our standardization for Q9, Q12 in the Free-viewing and Searching fields.
Algorithm standardization_Q2:
Input: batch_input (list of mixed numeric formats), GT_avg (default value if no number is found)
Output: cleaned_counts (list of integers)
1. Initialize an empty list cleaned_counts.
2. For each item in batch_input:
2.1 Attempt to directly convert the item to an integer:
a. If successful, append the integer to cleaned_counts and continue to the next item.
b. If conversion fails, proceed to step 2.2.
2.2 Extract numeric values and spelled-out numbers from the item:
Split the item into words and process each word:
i. Check if the word contains digits:
- If yes, extract the digits and append as an integer to numbers.
ii. Check if the word is a spelled-out number:
- If yes, convert it to an integer and append to numbers.
2.3 If no numbers were found, append GT_avg to cleaned_counts.
3. Return cleaned_counts.
Figure 9: Pseudo Code
Pseudo-code 9 shows our standardization for Q2 in the Subitizing field.
Algorithm standardization_Q1Q4Q5Q6Q7Q8Q10Q11Q13:
Input: batch_choice (list of entries with varied formats)
Output: cleaned_labels (list of extracted main labels)
1. Initialize an empty list cleaned_labels.
2. For each item in batch_choice:
2.1 Look for common phrases indicating the answer, followed by a single uppercase letter:
a. Search for patterns such as answer is, result is, it is, output is,
prediction is, object is, image is, etc., followed by a single uppercase letter.
b. If a match is found:
i. Extract the uppercase letter (label) from the matched pattern.
ii. Append the label to cleaned_labels.
c. If no match is found:
i. Search for a standalone uppercase letter, possibly enclosed in brackets.
ii. If found, append the letter to cleaned_labels.
iii. If no letter is found, append an empty string (“) to cleaned_labels.
3. Return cleaned_labels.
Figure 10: Pseudo Code
Pseudo-code 10 shows our standardization for Q1, Q4, Q5, Q6, Q7, Q8, Q10, Q11, Q13 in the Prominence, Prioritizing, Free-viewing, and Searching fields.
Algorithm standardization_Q3:
Input: batch_input (list of entries with varied formats)
Output: cleaned_labels (list of extracted and sorted labels)
1. Initialize an empty list cleaned_labels.
2. For each item in batch_input:
2.1 Search for specific patterns indicating a list of answers:
a. Look for phrases such as answer is, should be, order is, orders are, ranking is, etc., followed by a sequence of uppercase letters (possibly separated by spaces or symbols like
b. If a match is found:
i. Extract only the uppercase letters from the matched sequence.
ii. Append the extracted letters as a single string to cleaned_labels.
2.2 If no specific pattern is matched:
a. Search for all uppercase letters throughout the item.
b. Concatenate the found letters into a single string.
c. Append the concatenated string to cleaned_labels.
2.3 If no uppercase letters are found in the item, append an empty string (“).
3. Return cleaned_labels.
Figure 11: Pseudo Code
Pseudo-code C shows our standardization for Q3 in the Prioritizing field.
Appendix D Additional Qualitative Examples and Results
Fig. 12 and Fig. 13 show additional qualitative examples and results.
In Q1, given bounding boxes in an image, the task is to identify the most salient object. We observe that GPT-4o generates the incorrect answer (B), while LLaVA-OneVision produces unrelated text and selects an incorrect answer (E). Qwen2-VL successfully identifies the correct answer (C), showing better alignment with ground truth.
In Q2, the task is to count the number of salient objects in an image. The ground truth is 5. Qwen2-VL predicts six, and LLaVA-OneVision predicts four. Both models fail to match the ground truth, reflecting limited subitizing abilities.
In Q3, the task requires ranking the saliency of objects in order (e.g., ABC). The ground truth is ABC. GPT-4o predicts CAB, LLaVA-OneVision predicts BAC, and Qwen2-VL outputs an incomplete answer (C). None of the models produce the correct ranking.
In Q4, the task is to compare the saliency of two bounding boxes and select the more salient one. The ground truth is A. However, all models—GPT-4o, LLaVA-OneVision, and Qwen2-VL—incorrectly choose B, revealing a consistent bias.
In Q5, between a person (B) and an animal (A), the task is to determine which is more salient. The ground truth is B (the person). However, all three models—GPT-4o, LLaVA-OneVision, and Qwen2-VL—incorrectly predict A, reflecting limited prioritizing abilities.
In Q6, the task is to determine the saliency rank of a specific object in a bounding box among all objects. The ground truth is C (the third most salient). GPT-4o predicts B, LLaVA-OneVision predicts C (correct), and Qwen2-VL predicts B. Only LLaVA-OneVision aligns with the ground truth.
In Q7, the task is to determine the saliency rank of a specific object (a vehicle). The ground truth is A (the most salient). GPT-4o predicts B, LLaVA-OneVision predicts C, and Qwen2-VL predicts B. None of the models produce the correct ranking.
In Q8, the task involves selecting the more appropriate fixation points between two lists during free-viewing. The ground truth is B. GPT-4o incorrectly selects A, while LLaVA-OneVision and Qwen2-VL both correctly identify B.
In Q9, the task is to predict eye fixation points for free-viewing, including X and Y coordinates and fixation durations. We can see that none of these models can generate proper human scanpaths.
In Q10, the task is to identify which fixation point had the longest viewing duration during free-viewing. The ground truth answer is A (0.48, 0.50). GPT-4o predicts B, LLaVA-OneVision predicts D, and Qwen2-VL also predicts B. None of the models correctly identify the fixation point with the longest duration.
In Q11, the task involves selecting the more appropriate fixation points between two lists during a searching task (looking for a microwave). The ground truth answer is B. GPT-4o and Qwen2-VL incorrectly select A, while LLaVA-OneVision correctly selects B, aligning with the ground truth.
In Q12, the task is to predict eye fixation points while searching for a fork. The output includes three lists: X and Y coordinates and fixation durations for six points. Example outputs are provided for another image, but the document does not include detailed quantitative comparisons of the models’ predictions to ground truth, leaving their relative performance unclear.
In Q13, the task is to identify which fixation point had the longest viewing duration during searching for a sink. The ground truth answer is D (0.18, 0.59). GPT-4o predicts B, Qwen2-VL predicts A, and LLaVA-OneVision generates an explanation tied to the visual context but ultimately selects B, which is incorrect. None of the models produce the correct answer.
Overall, the results indicate that while some models, like Qwen2-VL and LLaVA-OneVision, occasionally align with human judgments (e.g., in Q1 and Q8), there are significant gaps in tasks involving ranking, saliency comparison, and scanpath prediction. Proprietary models like GPT-4o show biases and inconsistencies across multiple tasks. These results also demonstrate significant challenges for all models in accurately predicting fixation points and durations, particularly in tasks requiring nuanced alignment with human visual behavior. While LLaVA-OneVision shows occasional alignment (e.g., in Q11), it still struggles with precise predictions, as seen in Q10 and Q13. Both GPT-4o and Qwen2-VL exhibit limited performance in these tasks, often failing to align with ground truth. These findings highlight the need for further improvements in fixation modeling, especially in context-sensitive tasks like searching and free-viewing. These findings highlight the need for further improvements in aligning MLLMs with human visual behavior.
Figure 12: Additional Qualitative Examples and Results
Figure 13: Additional Qualitative Examples and Results
Appendix E Field-Specific Hints in each field
The Field-Specific Hints for each question type is as follows: {“Q1”: salient_hint, “Q2”: salient_hint, “Q3”: ranking_hint, “Q4”: ranking_hint, “Q5”: ranking_hint, “Q6”: ranking_hint, “Q7”: ranking_hint, “Q8”: fixation_hint, “Q9”: fixation_hint, “Q10”: fixation_hint, “Q11”: fixation_hint, “Q12”: fixation_hint, “Q13”: fixation_hint} .
And the full text for the hints:
salient_hint = “The detection of salient objects aims to simulate the human visual perception system by identifying and localizing the most visually striking object(s) in a scene [61]. Previous research suggests that the most salient object is the one that attracts the highest proportion of fixations, as indicated by the agreement between fixation patterns and saliency judgments [4]. In general, there are two primary priors: objects closer to the viewer are perceived as more salient, and salient objects often appear near the center of the scene [9]. Additionally, cues such as color contrast, spatial bias, and depth contrast also influence saliency [9].”
ranking_hint = “Ranking the saliency is to simulate the sequential shifting of human attention across objects during non-task-oriented image viewing, reflecting the limited capacity of the human visual system to process multiple visual inputs simultaneously [40]. In general, there are two primary priors: objects closer to the viewer are perceived as more salient, and salient objects often appear near the center of the scene. Additionally, cues such as color contrast, spatial bias, and depth contrast also influence saliency [9].”
fixation_hint = “Human fixations refer to the temporal sequence of locations in an image where individuals focus their gaze [22]. These fixations are typically recorded using an eye tracker under controlled laboratory conditions [22]. A scanpath includes not only the fixation locations but also the associated durations at each location [5]. Both low-level image properties and saliency, as well as high-level semantic information, serve as critical cues for predicting scanpaths [5]. The scanpath often begins at the center of the image.”
Appendix F Detailed Benchmark Results
[…]
Q1 Q2 Q2 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q9 Q10 Q11 Q12 Q12 Q13
[…]
       Baseline
Random guess 0.2806 0.1994 0.1909 1.9996 2.4336 0.0415 0.5 0.5 0.33 0.33 0.5 0.6079 0.8304 0.25 0.5 0.5425 0.8160 0.25
Proprietary MLLMs
GPT4-o [1] 0.3946 0.3139 0.4512 1.3445 3.2614 0.0924 0.6112 0.6260 0.3399 0.4725 0.5210 0.5917 0.8042 0.2368 0.5514 0.5106 0.7834 0.2731
Gemini-1.5-Flash [45] 0.3886 0.3323 0.5106 1.3070 3.0978 0.0427 0.5375 0.6855 0.3435 0.4560 0.5297 0.6128 0.8392 0.2417 0.5297 0.5199 0.8206 0.2963
OpenSource MLLMs
MiniGPT4-v2 [62] 0.1545 0.0879 0.1029 2.4939 3.0657 0.0000 0.3675 0.0968 0.2323 0.0495 0.1782 0.5017 0.4990 0.1874 0.3622 0.5364 0.5051 0.1157
DeepSeek-VL [34] 0.3655 0.3223 0.4544 1.2471 2.2514 0.0000 0.5750 0.6423 0.3667 0.4890 0.5149 0.5118 0.5450 0.1862 0.5514 0.4790 0.7130 0.1806
Idefics [26] 0.2488 0.1905 0.2386 2.1971 3.1163 0.0058 0.4638 0.4309 0.2482 0.0220 0.5050 0.5422 0.6236 0.1540 0.3946 0.5293 0.4895 0.1250
Idefics2 [26] 0.3067 0.2015 0.2990 2.2078 3.8573 0.0069 0.4950 0.6210 0.3301 0.4396 0.4926 0.5151 0.5525 0.2440 0.5297 0.5322 0.7619 0.2639
Idefics3 [25] 0.3552 0.2149 0.4852 1.1055 1.9567 0.0035 0.5625 0.5323 0.3839 0.4396 0.5062 0.5329 0.6310 0.1876 0.4541 0.5407 0.6856 0.1389
LLaVA-Next [32] 0.3460 0.3223 0.3961 1.3727 2.2675 0.0242 0.6000 0.6048 0.3460 0.4011 0.5025 0.5249 0.5998 0.1646 0.5351 0.5083 0.7883 0.1343
LLaVA-OneVision [28] 0.4035 0.4640 0.4517 1.1206 1.9671 0.0381 0.6737 0.5806 0.3362 0.3077 0.5099 0.5490 0.8177 0.2762 0.6432 0.5047 0.7295 0.3611
mPLUG-Owl3 [58] 0.3076 0.3309 0.2688 2.8941 4.6162 0.0000 0.6100 0.5565 0.3521 0.4231 0.4418 0.5028 0.4941 0.1908 0.4595 0.5368 0.5592 0.1019
Qwen2-VL [52] 0.4076 0.4103 0.5090 1.4698 2.6747 0.0589 0.6687 0.6016 0.4389 0.3791 0.5718 0.5199 0.7178 0.0828 0.6919 0.4610 0.8200 0.1898
InternVL2.0 [8] 0.3082 0.3101 0.3405 2.3458 3.6068 0.0473 0.5738 0.5806 0.3472 0.5165 0.4715 0.5028 0.7090 0.2195 0.4378 0.5016 0.7282 0.1759
MiniCPM-V 2.6 [57] 0.3476 0.3748 0.4491 1.2471 2.1137 0.0370 0.4812 0.6129 0.3447 0.3242 0.4728 0.5132 0.5589 0.1438 0.4757 0.4953 0.7107 0.2269
Table 11: HVSBench Leaderboard. The results of 13 leading MLLMs reveal significant room for improvement.
Table 11 shows the detailed benchmark on all question types. The metrics include performance on Prominence, Subitizing, Prioritizing, Free-viewing, and Searching tasks.
Analysis of Results
Prominence (Q1):
Task Summary: Determine the most salient object in an image. Analysis:
  • LLaVA-OneVision achieves the highest accuracy (46.4%), significantly outperforming all other models, indicating better alignment with human judgments of prominence.
  • Qwen2-VL (41.03%) follow, showing strong but less consistent performance.
  • Many open-source models, such as MiniGPT4-v2 (8.79%) and Idefics (19.05%), struggle, showing a lack of saliency understanding.
Subitizing (Q2):
Task Summary: Predict the number of salient objects in the image. Analysis:
  • Qwen2-VL and Gemini-1.5-Flash achieve the highest accuracy (51.06% and 50.90%), suggesting strong counting abilities.
  • Idefics3 excels in minimizing MAE (1.11) and RMSE (1.96), outperforming proprietary models like GPT4-o.
  • Open-source models like MiniGPT4-v2 and Idefics suffer high MAE and RMSE, indicating challenges in subitizing.
Prioritizing (Q3–Q7):
Task Summary: Rank or compare the saliency of objects or bounding boxes. Analysis:
  • Qwen2-VL and LLaVA-OneVision consistently rank highly across all prioritizing questions, with Qwen2-VL leading in Q3 accuracy (5.89%). We also observe the outperformance of DeepSeek-VL in Q5 and Q7.
  • Proprietary models, particularly GPT4-o, show competitive performance on Q3, indicating decent saliency ranking capabilities.
  • MiniGPT4-v2 and Idefics struggle with near-zero accuracy for Q3, showing a limited ability to rank saliency effectively.
Free-Viewing (Q8–Q10):
Task Summary: Predict or identify free-viewing scanpaths and their properties. Analysis:
  • Gemini-1.5-Flash leads in free-viewing accuracy (52.97%) and scanpath similarity metrics (M-Dir: 61.28%, M-Pos: 83.92%), demonstrating strong alignment with human-like scanpaths.
  • LLaVA-OneVision achieves competitive performance in scanpath metrics (M-Dir: 54.90%, M-Pos: 81.77%), slightly behind Gemini-1.5-Flash.
  • Most open-source models, like MiniGPT4-v2 and InternVL2.0, perform worse than random guessing, highlighting the difficulty of free-viewing tasks for these models.
Searching (Q11–Q13):
Task Summary: Predict or identify searching scanpaths and their properties. Analysis:
  • Qwen2-VL leads in searching accuracy (69.19%) and achieves the highest M-Dir (82.00%), indicating excellent alignment with human search behaviors.
  • LLaVA-OneVision is a strong second in accuracy (64.32%) but performs badly in scanpath similarity metrics (M-Dir: 50.47%, M-Pos: 72.95%).
  • Most models, including proprietary ones like GPT4-o, struggle with M-Dir and M-Pos metrics, showing limited scanpath prediction capabilities.
Overall Performance:
  • Best Performers: Qwen2-VL achieves the highest overall accuracy (40.76%), demonstrating robust performance across almost all tasks, particularly in Subitizing and Searching. LLaVA-OneVision closely follows (40.35%) and excels in Prominence and Free-viewing tasks. Gemini-1.5-Flash outperforms on Free-viewing and Subitizing tasks but lags slightly in overall accuracy.
  • Proprietary vs. Open-Source: Proprietary models like GPT4-o remain strong in specific tasks like Prominence but are often surpassed by advanced open-source models like Qwen2-VL and LLaVA-OneVision. Most models (e.g., MiniGPT4-v2, Idefics) underperform across most tasks, indicating challenges in aligning with human visual system behaviors.
Key Observations:
  • Scanpath Prediction Gaps: Most models, including top performers, fail to achieve human-like scanpath similarity in M-Dir and M-Pos, even underperforming random guessing in Free-viewing tasks.
  • Task Variability: Searching tasks are relatively easier for MLLMs, as they involve structured objectives, while Free-viewing is the most challenging due to its reliance on unconstrained, human-like behavior.
  • Room for Improvement: Despite some standout performances, all models exhibit significant gaps in HVS alignment, emphasizing the need for better modeling of saliency, scanpaths, and human-like visual prioritization.
F.1 Ablation on Model Size
We evaluate the impact of model size on the performance in HVSBench by testing models with different parameter counts. For this study, we select two representative methods: DeepSeek-VL [34] and GPT-4o [1], from both open-source and proprietary MLLMs, to provide a comprehensive analysis. As shown in Table 4, larger MLLMs generally outperform smaller ones across all metrics. It suggests that increasing model size leads to better alignment with HVS for MLLMs. Refer to supplementary for more experiments.
[…]
  GPT4-o mini [1] N/A 0.3126 0.4480 0.3312 0.3560 0.3766
GPT4-o [1] N/A 0.3139 0.4512 0.3621 0.3737 0.4015
DeepSeek-VL [34] 1.3B 0.1758 0.2513 0.2950 0.3188 0.2843
7B 0.3223 0.4544 0.3327 0.3445 0.3516
Qwen2-VL [52] 2B 0.0220 0.3499 0.1100 0.2439 0.2643
7B 0.4103 0.5090 0.3901 0.3182 0.4214
72B 0.4957 0.4889 0.4158 0.4299 0.5810
Table 12: Ablation study of The number of params. PO, SU, PI, FV and SE means “Prominence”, “Subitizing”, “Prioritizing”, “Free-viewing”, “Searching”, respectively.
F.2 Discussion: Why do MLLMs work on HVS-related tasks?
The design of MLLMs (particularly transformer-based architectures) allows approximations of human visual capabilities. For Prominence, Subitizing and Prioritizing, [44] leverages MLLMs to derive visual saliency hierarchies as the guidance for saliency prediction, proving MLLMs’ potential capacity to mimic human visual prioritization. For Free-Viewing and Searching, [56] can predict scanpaths by simulating gaze patterns using Transformer-based attention mechanisms. For Prominence, Subitizing and Prioritizing, [47] shows that human-like saliency can be simulated by attention layers inherently learning to weight salient regions of an input image, mirroring human prioritization of significant elements. We highlight the value of benchmarking HVS alignment, as it benefits many tasks.
Appendix G More Related Work
Human Visual System (HVS). The HVS has long been studied for its unique ability to process visual information efficiently and selectively. Computational modeling of HVS has gained significant traction in the fields of computer vision and cognitive neuroscience, aiming to replicate human-like attention and perception in artificial systems. [20] shows how visual saliency guides human gaze patterns. Recent advancements in deep learning have incorporated human attention models into computer vision tasks, enabling better predictions of free-view human gaze [24, 11]. These approaches provide insight into how human cognition hierarchically processes visual information. HVS also demonstrates sequential and temporal fixation patterns, critical for understanding complex scenes [17]. The study of the HVS has led to significant improvements and inspired new models in machine learning, such as attention models [51]. It is crucial to conduct further research into the HVS due to its potential to advance the development of AGI.
Multimodal Large Language Models (MLLMs). MLLMs [2] have emerged as a significant advancement in artificial intelligence, extending the capabilities of large language models to process and reason about both visual and textual information. By utilizing the open-source LLM [48, 49, 46, 59] and the key idea of constructing visual instruction data, some powerful MLLMs have been proposed such as LLaVA [32] and MiniGPT-4 [62]. These models have shown their ability in general visual tasks.
Despite these advancements, questions about how MLLMs perceive and process visual information remain largely unexplored. It is unclear whether MLLMs fixate on regions of interest similar to humans or follow a comparable temporal sequence when perceiving images. Furthermore, further research in this area is hindered by the absence of standardized evaluation protocols and benchmarks.